Fermi National Laboratory
 
Operation Summary for November 14, 2011 - November 16, 2011
 

Wednesday November 16

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the MiniBooNE, NuMI, and the Meson FTBFT-978 experiments taking beam.

During the day shift, a Muon person began a study at 8 AM. Operators reported at 8:17 AM that a Booster RF station (BRF19) had tripped off several times in the last few minutes; RF personnel investigated. Operators reported at 9:23 AM that a Controls person changed a BPM crate at MI-30. Operators search and secure MTA and Operations established beam to MTA at 10:37 AM. At 2:01 PM, a MTA power supply’s (E:UHB7) output dropped to zero; Power Supply personnel investigated. The Muon study ended at 3:58 PM.

During the evening shift, Operations disabled T-978 beam at 7:04 PM. Operators completed a tune of the 750 KeV line at 11:56 PM.

Thursday November 17

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 12:37 AM, operators reported that NuMI had tripped off twice due to high beam intensity; operators lowered the intensity to NuMI. Operators reported at 2:29 AM that a NuMI BPM (NBPM6N) stopped polling; it reset with a reboot. Operations established beam to T-978 at 7:34 AM. At 7:45 AM, Neutron Therapy Facility (NTF) personnel began preparing to treat a patient later in the morning.

During the day shift, at 8:42 AM, operators reported that BRF19 was tripping off with no beam; operators contacted RF personnel. Operators reported at 8:45 AM that twice during this shift a NuMI BPM system (NBPMOV) had taken down the beam permit; an Instrumentation person investigated. At 1:48 PM, a Linac quadrupole magnet power supply (L:QPS108) tripped off for the first of two times in nine minutes; it reset.

During the evening shift, at 4:46 PM, the NuMI beam permit tripped off when the output current of two power supplies drifted low; everything reset. At 4:47 PM, the NuMI beam permit tripped off for the first of two times in about two hours when a NBPMOV stopped polling; everything reset when rebooted. Operations disabled the T-978 beam at 7:10 PM. Operators reported short downtimes due to LRF3 and BRF1.

Friday November 18

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 3:31 AM, NBPMOV stopped polling and the remote reboot would not work; operators called in an Instrumentation person. Operations continued to deliver beam to NuMI. Operators turned on the Muon power supplies at 5:07 AM. Operations turned off the NuMI beam at 5:21 AM per instructions from a NuMI person due to several problems, one of which included a bad VME crate. Technicians replaced the NuMI VME crate and Operations resumed sending beam to NuMI at 6:48 AM. A Muon person began a study at 7 AM. Operations established beam to T-978 at 7:15 AM. At 7:50 AM, NTF personnel began preparing to treat a patient later in the morning.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The OpsCo approved of the plans by Booster, MI, and Muon personnel to conduct studies today. Operations will deliver protons to the experiments for the weekend.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/18/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory